OpenStudy (demonphannie):

What is the value of 3 to the power 2 over 3 to the power 4? 1 over 81 1 over 27 1 over 9 1 over 3 I WANT TO LEARN IT PLEASE EXPLAIN

OpenStudy (saadsword):

(3^2)/(3^4)=9/81=1/9

OpenStudy (saadsword):

choice 3

jhonyy9 (jhonyy9):

use formula a^x ---- = a^(x-y) a^y